OpenStudy (anonymous):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day? A. x = −9, x = 21 B. x = −6, x = 18 C. x = 6, x = −18 D. No solution

pooja195 (pooja195):

Add -14 to both sides. 2(|x−6|)+14+−14=38+−14 2(|x−6|)=24 Step 2: Divide both sides by 2. 2(|x−6|)/2=24/2 |x−6|=12 Now set up two equations one postitve one negative (sol)

pooja195 (pooja195):

x−6=12 x−6=-12 now solve for x

OpenStudy (anonymous):

so B is correct?

pooja195 (pooja195):

yep :)

OpenStudy (anonymous):

thnx guys :)